Use the Trapezoidal Rule, the Midpoint Rule, and Simpson's Rule to approximate the given integral with the specified value of n. (Round your answers to six decimal places.

\int_{\0 }^{pi/2} 3sqrt(1 + cos(x)) dx, n = 4
Answer #1

12 3 sqft (it cos (80).dy, m=4 1 Trapezondal Rule 3 Jul(a) +1 .da. n=4. The bapezoidal enle states that floo. dx x dx (flaco)6 2-055253 4 35193) = 24/59) - froy) = feb) = f(3) = Imally just eum wpabove valus and mus by mustry 16 =D [ 4.242640 +8.3222Noeleze finchen at there experts frete) - [ Deng) - () = 3tering til /2+) == poje = 0 ) - + (89) 32 Lim (20) 2, = US59054 $Date 2 fx2 + . ut affen-2) + 4 fron) + fran) where ax=b-a n & Ax=0 Duide interval (0,1/2] mto n=4 subhtervals of length are=1